Jefferson Lab HALL A Logbook for September 00 - All EXCEPT Automatic Entries

  • Main INDEX

    Saturday

  • 48786 09/30/00 22:42 Templon fake daq problems
  • 48775 09/30/00 21:52 Sawatzky beam trip mid-run
  • 48773 09/30/00 21:43 Templon detector to-do list
  • 48756 09/30/00 20:22 Sawatzky Run 1184 [EP]mom=0.789 not 0.805
  • 48744 09/30/00 19:18 Templon hana comments IV
  • 48743 09/30/00 19:14 Templon hana comments III
  • 48742 09/30/00 19:10 Templon daq problems
  • 48736 09/30/00 18:39 Templon CODA stopped counting
  • 48727 09/30/00 18:07 Templon hana comments resumed on HP machine
  • 48721 09/30/00 17:21 Templon halog screen grab on linux is BROKEN
  • 48719 09/30/00 17:12 Templon hana analysis notes
  • 48716 09/30/00 16:48 Templon deadtime, prescale factor change
  • 48701 09/30/00 14:55 LeRose Reboot DAQ
  • 48676 09/30/00 11:17 R. Michaels checks of event tag - looks good
  • 48645 09/30/00 07:42 K. McCormick magnet crash
  • 48625 09/30/00 04:21 Templon harp scan / bcm cal
  • 48617 09/30/00 03:52 Templon beam posn scan info
  • 48613 09/30/00 03:37 Templon optics target
  • 48588 09/30/00 01:57 R. Michaels kcoda does restart ET
  • 48586 09/30/00 01:40 Templon daq problem now really solved
  • 48580 09/30/00 01:07 Templon possible solution to daq problem
  • 48572 09/30/00 00:33 Templon Harp scans
  • Friday

  • 48566 09/29/00 23:25 Adaq daq problem info
  • 48563 09/29/00 23:21 Templon harps happy again
  • 48560 09/29/00 22:39 Templon coda not happy
  • 48559 09/29/00 22:38 Templon controlled access
  • 48555 09/29/00 22:04 Templon HARP scans not working
  • 48553 09/29/00 21:31 R. Michaels datamon fixed
  • 48548 09/29/00 20:40 Sawatzky pulser check of proton momentum trigger settings
  • 48545 09/29/00 20:37 Templon bcm calibration completed
  • 48544 09/29/00 20:28 Templon harp scan prep, hac inconvenience
  • 48539 09/29/00 20:05 Templon bcm calibration
  • 48537 09/29/00 19:59 Templon gas cerenkov HV turned on
  • 48518 09/29/00 18:10 Templon daq
  • 48515 09/29/00 17:56 Jack Crate resets.
  • 48513 09/29/00 17:47 Jack Left arm gas cherenkov.
  • 48512 09/29/00 17:42 Templon no data
  • 48497 09/29/00 16:43 R. Michaels DAQ/Trigger status
  • 48470 09/29/00 15:01 penel VDC HV reset
  • 48469 09/29/00 14:51 Templon various analysis notes
  • 48461 09/29/00 11:36 M Jones Problems with gas
  • 48460 09/29/00 11:34 M Jones changes in harp software
  • 48454 09/29/00 10:40 M Jones VDC top on Right HRS
  • 48453 09/29/00 10:20 Templon no trigger 3's
  • Thursday

  • 48446 09/28/00 17:55 templon cabling changes
  • 48432 09/28/00 12:08 Templon VDC cabling
  • 48430 09/28/00 09:33 Templon argon gas bottle change
  • Wednesday

  • 48429 09/27/00 21:17 R. Michaels DAQ/Trigger changes, Aug-Sep
  • Tuesday

    
    

    Monday